In 3 years time a pet mouse will be as old as his owner was 4 years ago. Their present total age is 13 years. Find the age of each now
Answer by BreeCure(1)

Let y be the age of the owner
let x be the age of the pet
Therefore: x+y=13
x+3=y-4
then solve for x
x+y=13
x=13-y
x+3=13-x-4
2x=6
x=3
then solve y
y=13-3
y=10
therefore the pet's age is 3years and the owner's age is 10years